the uncertainity in momentum of an electron is 1×10 -5 kg.m/s. the uncertainity in its position will be[h=6.62×10 -34 kg.m/s

  • the uncertainity in momentum of an electron is 1×10-5kg.m/s. the uncertainity in its position will be[h=6.62×10-34kg.m/s

Grade:11

1 Answers

Aarti Gupta
9 years ago
Given - Uncertainity in momentum , delta p = 1*10-5 kgm/s
Planck’s constant h = 6.62 * 10-34 kgm/s
Uncertainity in position,delta x = ?
Sol.-
According to Heisenberg uncertainity principle,
deltax * deltap = h/4pi
deltax = h/4pi * deltap
= 6.62 * 10-34/4*3.14*1*10-5
deltax = 5.27 * 10-30 m
Thus uncertainity in position = 5.27 * 10-30 m